OpenStudy (anonymous):

Two consecutive integers have a product of 6, what are the integers?

OpenStudy (anonymous):

let the two concecutive integers be x and x+1 as per question: x(x+1)=6 solve for x

OpenStudy (anonymous):

2(2+1)=6? ... 2 times 3 = 6?

OpenStudy (anonymous):

ya its right.........but we take it using a general no....

OpenStudy (anonymous):

So the final answer is 2,3?

OpenStudy (unklerhaukus):

but there are two sets of solutions

OpenStudy (unklerhaukus):

let the two consecutive integres be \( \{n,n+1\}\)\[(n)(n+1)=6\] \[n^2+n=6\]\[n^2+n-6=0\] \[(n-2)(n+3)=0\] \[n=2,-3\] so \[ \{n,n+1\}=\{2,2+1\}=\{2,3\}\]or\[\{n,n+1\}=\{-3,-3+1\}=\{-3,-2\}\]